OpenStudy (anonymous):

sin40 * cos20 + cos40 * sin20 help... where do I start!?

OpenStudy (ryan123345):

Hey there! I see that you are new to openstudy. This is a great website on getting help with questions of any subject. It is a fun way of learning! I wish you good luck on openstudy and dont forget to read the http://openstudy.com/code-of-conduct so you know the rules and regulations :) Hope you have fun!

OpenStudy (anonymous):

yes, wish i would have found this site earlier this semester! thanks!

OpenStudy (anonymous):

Recall the angle sum-difference identity for sine: \[\sin(\alpha+\beta)=\sin\alpha\cos\beta+\sin\beta\cos\alpha\] \[\sin40\cos20+\cos40\sin20=\sin(40+20)=\sin60\] (angles in degrees)